OpenStudy (anonymous):

]Identify the correct x intercept and y intercept of the equation x + 7y = -21. x intercept = (3, 0); y intercept = (0, 21) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (21, 0); y intercept = (0, 3) x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (anonymous):

y=mx+b y=(x/7)-21 b = -21 which equals y intercept

OpenStudy (lgbasallote):

x intercept..let y = 0 then solve for x x + 7y = -21 when y = 0 x + 7(0) = -21 solve for x y intercept...let x = 0 then solve for y x + 7y = -21 when x = 0 0 + 7y = -21 solve for y @geo_mitch you forgot something ;)

OpenStudy (lgbasallote):

got it @Megan10118765 ?

OpenStudy (anonymous):

Thank you for picking that up Igbasallote